A motorcycle is moving at 30m/s when the rider applies the brakes, giving the motorcycle a constant deceleration. During the 3.0s interval immediately after braking begins, the speed decreases to 15m/s.

(a) What distance does the motorcycle travel from the instant braking begins until the motorcycle stops?

(b) Sketch graphs of x, v, and a vs t. Explain the graphs.
Answer #1

inidial velocity (v.) = 30m/s tinal velocity (ro) - 15 m/s t=35 - V 3 at -u-ft t = f f = 30-15 - supe the distance inotorcyclS